OpenStudy (anonymous):

Factor completely: 4x2 + 25x + 6 (4x + 1)(x + 6) (4x + 6)(x + 1) (2x + 3)(2x + 2) (2x + 6)(2x + 1) @johnweldon1993

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I forgot how...>_<

OpenStudy (whpalmer4):

You don't know how to multiply \[(4x+1)(x+6)\]?

OpenStudy (mathstudent55):

This is how. Use FOIL: (a + b)(c + d) = ac + ad + bc + bd

OpenStudy (anonymous):

Ohh, I forgot! Thanks @mathstudent55

OpenStudy (anonymous):

How would I multiply unlike terms again? I forgot. (4x x 6)

OpenStudy (whpalmer4):

4x * 6 = 24x

OpenStudy (whpalmer4):

it's just 4x + 4x + 4x + 4x + 4x + 4x, after all...

OpenStudy (anonymous):

Oh, okay, I thought it was different. Guess not. > . <

OpenStudy (whpalmer4):

you can't add them...maybe that's what you were thinking of...

OpenStudy (anonymous):

Oh yeah, that's it. Can't add em. So now i have \[(4x^2+24x)(x+6)\]

OpenStudy (whpalmer4):

mmm.no....are you doing the first answer choice? \((4x+1)(x+6) = 4x(x+6) + 1(x+6)\) by the distributive property of multiplication \[4x(x+6)+1(x+6) = 4x*x + 4x*6 + 1*x + 1*6\]again, distributive property collecting like terms, we get\[4x^2+24x+1x+6=4x^2+25x+6\]

OpenStudy (whpalmer4):

basically, each piece of the first thing multiplies each piece of the second thing, and then you collect it all together...

OpenStudy (anonymous):

Ohhh. I get it. Thanks.

OpenStudy (whpalmer4):

It might be easier to see with a numeric example. If we multiply 12 * 25, that's the same as \[(10+2)*(20+5) = 10*20 + 10*5 + 2*20 + 2*5\]\[ = 200 + 50 + 40 + 10 = 300\]